1. Suppose that the Matrix Corporation has $120 million of assets, all equity financed, and that the firm has 6 million shares of stock outstanding valued at $20 per share. Now that management has identified investment opportunities requiring $60 million of NEW funds and it can be raised in one of the following 3 ways:

• Strategy 1: Issue $60 million equity.

• Strategy 2: Issue $30 million of equity and borrow $30 million with r = 8%.

• Strategy 3: Borrow $60 million with r = 8%.

a) Suppose the firm’s rate of return is r ∗ = 15%, compute the earnings per share for the 3 financing strategies listed above.

b) Suppose the firm’s rate of return is r ∗ = 10%, compute the earnings per share for the 3 financing strategies.

c) Suppose the firm’s rate of return is r ∗ = 5%, compute the earnings per share for the 3 financing strategies.
Answer #1

As per the given problem,

Requirement of new funds = $60 million.

Share value per share = $20

No.of new shares issued under 3 strategies:

Particulars Strategy 1 Strategy 2 Strategy 3
New funds raised $60 million $60 million $60 million
New funds raised through equity (a) $60 million $30 million -
Share price(b) $20 $20 $20
No.of shares issued (c)= a/b 3 million shares 1.5 million shares 0
Existing shares (d) 6 million shares 6 million shares 6 million shares
Total equity shares (e)= c+d 9 million shares 7.5 million shares 6 million shares

Rate of interest for borrowings = 8%

Case 1: Firms rate of return (r) = 15%

Particulars strategy 1-100% Equity strategy 2- 50%Equity,50%debt Strategy 3-100% debt
New funds $60 million $60 million $60 million
Old funds $120 million $120 million $120 million
Total funds(a) $180 million $180 million $180 million
Rate of return (b) 15% 15% 15%
Earnings before interest & taxes(EBIT) c=(a*b) $27,000,000 $27,000,000 $27,000,000
(-) Interest (Debt * 8%) (d) -

($2,400,000)

($60 million * 8%)

($4,800,000)

($60 million * 8%)

Earnings before taxes (e=c-d) $27,000,000 $24,600,000 $22,200,000
(-) Taxes (f) - - -
Earnings after taxes (g)=(e-f) $27,000,000 $24,600,000 $22,200,000
No.of equity shares (h) 9,000,000 shares 7,500,000 shares 6,000,000 shares

Earnings per share

(i)=(g) /( h)

$3 per share $3.28 per share $3.7 per share

Case 2: Firms rate of return (r) = 10%

Particulars strategy 1-100% Equity strategy 2- 50%Equity,50%debt Strategy 3-100% debt
New funds $60 million $60 million $60 million
Old funds $120 million $120 million $120 million
Total funds(a) $180 million $180 million $180 million
Rate of return (b) 10% 10% 10%
Earnings before interest & taxes(EBIT) c=(a*b) $18,000,000 $18,000,000 $18,000,000
(-) Interest (Debt * 8%) (d) -

($2,400,000)

($60 million * 8%)

($4,800,000)

($60 million * 8%)

Earnings before taxes (e=c-d) $18,000,000 $15,600,000 $13,200,000
(-) Taxes (f) - - -
Earnings after taxes (g)=(e-f) $18,000,000 $15,600,000 $13,200,000
No.of equity shares (h) 9,000,000 shares 7,500,000 shares 6,000,000 shares

Earnings per share

(i)=(g) /( h)

$2 per share $2.08 per share $2.2 per share

Case 3: Firms rate of return (r) = 5%

Particulars strategy 1-100% Equity strategy 2- 50%Equity,50%debt Strategy 3-100% debt
New funds $60 million $60 million $60 million
Old funds $120 million $120 million $120 million
Total funds(a) $180 million $180 million $180 million
Rate of return (b) 5% 5% 5%
Earnings before interest & taxes(EBIT) c=(a*b) $9,000,000 $9,000,000 $9,000,000
(-) Interest (Debt * 8%) (d) -

($2,400,000)

($60 million * 8%)

($4,800,000)

($60 million * 8%)

Earnings before taxes (e=c-d) $9,000,000 $6,600,000 $4,200,000
(-) Taxes (f) - - -
Earnings after taxes (g)=(e-f) $9,000,000 $6,600,000 $4,200,000
No.of equity shares (h) 9,000,000 shares 7,500,000 shares 6,000,000 shares

Earnings per share

(i)=(g) /( h)

$1 per share $0.88 per share $0.7 per share